OpenStudy (anonymous):

HELP! **PROOF I GIVE MEDALS :) Given:In triangle ABC, ∠ A=135º Prove:∠B≠45º(write an indirect proof)

OpenStudy (anonymous):



OpenStudy (raden):

actually, <C is given :)

OpenStudy (anonymous):

is c 45

OpenStudy (raden):

depends in ur question but C cant be 45, because <A = 135 A+C = 180, it is not true because B = 0

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

if b is not 45 then it must be >0 and <45

OpenStudy (raden):

ah, that is the answer :) prove by using contradiction, assuming be right B = 45 degrees we knowed tha the sum interior angles in triangle = 180 A + B + C = 180 135 + 45 + C = 180 180 + C = 180 C = 0

OpenStudy (raden):

if C = 0 then it is not a triangle so, proof that C cant be 0

OpenStudy (raden):

in other words, in a triangle if the measure <A = 135 then ∠B≠45º

OpenStudy (anonymous):

and c>0

OpenStudy (anonymous):

oh ok sory I couldn't answer b4 myinternect went topellet

OpenStudy (anonymous):

and c must not be 45